# Decoupled Data Consistency with Diffusion Purification for Image Restoration

### v5

@@ 2.3 Diffusion Purification
Recently, diffusion purification has shown to be an effective approach for removing adversarial perturbations [22, 31] and for image editing [32]. Specifically, for a given noisy or perturbed image $\bm{x}_{\textrm{adv}}$, this process involves first running the diffusion forward process as described in Equation (2) starting from $\bm{x}_{\textrm{adv}}$ to some intermediate timestep $T$. In the discrete case, $\bm{x}_{T}$ takes the form
$\bm{x}_{T}=\sqrt{\bar{\alpha}_{T}}\bm{x}_{\textrm{adv}}+\sqrt{(1-\bar{\alpha}_{T})}\bm{\epsilon}\>.$ (5)
@@
Starting with $\bm{x}_{T}$, we implement the reverse diffusion process with Equation (3) to produce a “purified” image $\bm{x}_{\textrm{purified}}$, effectively eliminating the adversarial perturbations. The rationale behind this is that by carefully selecting $T$, the adversarial noise within $\bm{x}_{\textrm{adv}}$ becomes overwhelmed by the noise introduced during the forward phase, yet the core semantic information is preserved. Consequently, executing the reverse diffusion process leads to a “clean” image reconstruction. In our work, we use diffusion purification as the backbone of our approach, yet with a distinct objective diverging from existing methods that use diffusion purification to enhance the robustness of DNN-based classifiers [22] or remove artifacts in medical imaging [33, 31].
3 Proposed Methods
In this section, we first introduce our method for image restoration based upon variable splitting (Section 3.1). Then, in Section 3.2, we discuss enforcing data consistency, followed by our proposed method to use diffusion models as an image prior via diffusion purification (Section 3.3). The entire procedure of the proposed method is outlined in Algorithm 1.
3.1 Methods with Variable Splitting
@@
Our proposed algorithm is inspired by the classical variable splitting technique [34, 35, 36, 37] in optimization. Specifically, consider the following regularized optimization formulation for the image restoration problem in Equation 1,
$\min_{\bm{x}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{x}),$
where $\mathcal{R}(\cdot)$ is the regularization for image prior, and $\lambda>0$ controls the penalty strength. The idea of variable splitting is to introduce an auxiliary variable $\bm{v}=\bm{x}$ to pose an equivalent problem
$\min_{\bm{x},\bm{v}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{v}),\;\mathrm{s.t.}\;\bm{x}=\bm{v}.$ (6)
@@
This constrained optimization problem can be relaxed into an unconstrained problem by the half quadratic splitting technique (HQS) [38, 37], which we can reformulate as
$\min_{\bm{x},\bm{v}}\underbrace{\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\bm{v}\|_{2}^{2}+\lambda\mathcal{R}(\bm{v})\>,}_{=:\mathcal{L}(\bm{x},\bm{v})}$ (7)
@@
for some properly chosen $\mu>0$. Then, to solve (7), we can use an iterative method such that for every iteration $k$, we can leverage the extra variable $\bm{v}$ to split the optimization into two sub-problems of the form
$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\mathcal{L}(\bm{x},\bm{v}_{k-1}),$ (8)
$\bm{v}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{v}}\;\mathcal{L}(\bm{x}_{k},\bm{v}),$ (9)
that decouples the data consistency term $\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}$ from the regularization term $\mathcal{R}(\bm{v})$. Then, instead of having an explicit regularizer $\mathcal{R(\cdot)}$ (e.g., $\ell_{1}$-norm to induce sparsity), we can use implicit regularizers. More specifically, we can constrain the output space as the range of the diffusion model via diffusion purification that acts as a form of regularization or prior. To enforce this implicit regularizer, our method involves alternating between the two following steps:
  1. 1.
    Data Consistency. Generate measurement-consistent reconstructions, which may have various artifacts, by solving the data consistency optimization problem Equation 8.
  2. 2.
    Regularization via Diffusion Purification. Refine the obtained reconstruction from the previous step with a pre-trained diffusion model via diffusion purification.
In the following subsections, we dive into the technical details of each step and present an illustration of our method in Figure 2. We also conduct an ablation study in Section 5.3, demonstrating that alternating between these two steps can decrease the distance between the reconstructed and the target image.
Algorithm 1 Image Restoration with Decoupled Data Consistency and Diffusion Model Prior
@@
measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained model $\bm{s}_{\bm{\theta}}$, total number of iterations $K$, decaying time schedule $\{T_{k}\}_{k\in[K]}$ for diffusion purification, iteration number $\tau$, and $\bm{v}_{0}=\bm{0}\in\mathbb{R}^{n}$
@@
     $\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau)$triangleright solve via gradient based methods
@@
3.2 Enforcing Data Consistency
Recall that the sub-problem in Equation (8) for maintaining data consistency can be written as
$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}+\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}\>.$
The difficulty behind directly solving the above problem lies in selecting the proper hyperparameter $\mu$, as it may need to be frequently tuned. To tackle this issue, we instead seek an approximate solution $\bm{x}_{k}$ that minimizes solely the data fidelity term $\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$ but starting the optimization with $\bm{v}_{k-1}$ as the initial point and run a limited number of $\tau$ gradient steps, resulting in a measurement-consistent reconstruction $\bm{x}_{k}$. The rationale behind this is two-fold: (i) by taking $\bm{v}_{k-1}$ as the initial point, we can obtain an output $\bm{x}_{k}$ that is sufficiently close to $\bm{v}_{k-1}$ without having to tune $\mu$, and (ii) taking a limited number of gradient steps $\tau$ reduces the computational burden. Overall, for simplicity, we denote this step as
Lastly, we want to highlight that our technique for enforcing data consistency is significantly more efficient than existing methods, as our method does not require back propagation over the entire score function (i.e., the neural network) and hence is (very) computationally efficient. Furthermore, to enforce data consistency, we can use any momentum-based gradient descent optimizer (e.g., SGD, ADAM) for Equation 10. Throughout the experiment we fix the number of gradient steps $\tau$ for each round of the data fidelity optimization. When the measurements are corrupted by noise, one can further reduce the number of iteration to prevent overfitting. We discuss our choice of optimizers in Appendix A.
3.3 Implicit Regularization by Diffusion Purification
Due to the underdetermined nature of the problem (1), the initial reconstruction $\bm{x}_{k}$ obtained from Equation 10 will suffer from task specific artifacts, as shown in Figure 3 (the third column). The next step is to refine $\bm{x}_{k}$ via enforcing the image prior learned by the diffusion model. To do so, we propose to utilize diffusion purification described in Section 2.3 as an implicit regularizer for effectively removing the artifacts in $\bm{x}_{k}$ as shown in Figure 3. More specifically, let $T_{k}$ denote the diffusion purification strength at iteration $k$. Following the diffusion purification step in Equation (5), we first define $\bm{x}_{k,0}:=\bm{x}_{k}$ and submerge this sample with noise up to time $T_{k}$:
$\bm{x}_{k,T_{k}}=\sqrt{\bar{\alpha}_{T_{k}}}\bm{x}_{k,0}+\sqrt{1-\bar{\alpha}_{T_{k}}}\bm{\epsilon}.$ (11)
Then, we run unconditional reverse sampling via the reverse SDE in Equation 3 starting from $\bm{x}_{k,T_{k}}$ to obtain the output $\bm{v}_{k}$. Concretely, the discretized version of the reverse SDE can be described as
$\bm{x}_{k,t-1}=$ $\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{k,t}+\beta_{t}\bm{s}_{\bm{\theta}}(\bm{x}_{k,t},t)\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$ (12)
for $t=T_{k},T_{k}-1,\ldots,1$. Here, recall that $\bm{s}_{\bm{\theta}}(\bm{x},t)$ denotes the score function which can be estimated by a pre-trained unconditional diffusion model. For simplicity, we denote this overall regularization step by
$\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},T_{k}).$ (13)
For the choice of $T_{k}$, we propose to use a linear decaying schedule for robust performance, which is capable of addressing the “hallucination” issue brought by the stochasticity of the diffusion purification process. In the early iterations of our method, when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though may be inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data consistency optimization. As $T_{k}$ keeps decreasing, the algorithm will converge to an equilibrium point which is both consistent with the measurements and free from the artifacts. More technical details are deferred to Appendix A, and we study the convergence in Section 5.3.
4 Adaptability to Accelerated Samplers & Latent Diffusion
One of the main drawbacks of diffusion models is the increased inference time of the iterative sampling process compared to traditional methods such as variational autoencoders (VAEs) [39] and generative adversarial networks (GANs) [40]. In the literature, many advancements have been made to address this issue, which involve either decreasing the number of iterations via accelerated sampling techniques such as DDIMs [11] or carrying out the sampling process in a lower-dimensional, latent space [24, 25]. In the following sections, we briefly introduce these techniques and demonstrate how they can be used to decrease the inference time of diffusion purification, thereby accelerating our algorithm.
4.1 Faster Diffusion Purification
@@
To improve sampling efficiency, we can adopt the use of accelerated samplers to decrease the sampling time of the reverse process of diffusion purification. For example, we can use the accelerated sampler DDIM, which constructs a non-Markovian chain of the reverse process. By doing so, we can perform each round of diffusion purification with as few as $20$ unconditional reverse sampling steps (i.e., $20$ number of function evaluations (NFEs)), with minimal performance degradation.
Here, we would like to remark that existing methods such as DPS do not have this advantage. DPS typically requires approximately $1000$ reverse sampling steps, as it requires more iterations to ensure “sufficient” data consistency steps.
In the following, we discuss the adaptability of our approach using one-step sampling approaches, including consistency models and approximations via Tweedie’s formula.
Consistency Models.
Recently, Song et al. [19] proposed a class of diffusion models termed consistency models (CMs), which can effectively generate high-quality samples in a single (or at least very few) time steps. CMs build on the fact that there exists a deterministic ODE for every corresponding SDE sampling trajectory of Equation 3 in the form of
with which it shares the same probability densities $p_{t}(\bm{x}_{t})$ for all $t\in[0,1]$ [9]. This implies that for any image $\bm{x}_{0}$, there exists a unique ODE trajectory $\{\bm{x}_{t}\}_{t=0}^{1}$. By leveraging this property, CMs train a deep network $f_{\bm{\theta}}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ parameterized by $\bm{\theta}$ to map any point $\bm{x}_{t}$ to its corresponding $\bm{x}_{0}$ in a single step.
This approach can be seamlessly incorporated into our framework. Specifically, with a pre-trained consistency model $f_{\bm{\theta}}$, the diffusion purification step in Equation 13 in Algorithm 1 can be performed in a single step as
$\bm{v}_{k}=\texttt{DPUR}(\bm{f}_{\bm{\theta}},\bm{x}_{k},T_{k}):=f_{\bm{\theta}}(\bm{x}_{k},T_{k}).$ (15)
Approximation via Tweedie’s Formula.
Under a Gaussian assumption, Tweedie’s formula [29, 30] provides a way of computing the posterior mean by using
$\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})),$
@@
where $\bm{x}_{0}$ denotes a sample from the “clean” data distribution. Equipped with this formula, instead of running the reverse SDE in Equation 12 for diffusion purification, one can estimate the ground truth in a single step using this formula. Specifically, we can approximate $\bm{v}_{k}$ using
$\bm{v}_{k}$ $=\mathbb{E}\left[\bm{x}_{k,0}\mid\bm{x}_{k,T_{k}}\right]\approx\frac{1}{\sqrt{\bar{\alpha}_{T_{k}}}}(\bm{x}_{k,T_{k}}+(1-\bar{\alpha}_{T_{k}})\bm{s}_{\bm{\theta}}(\bm{x}_{k,T_{k}},T_{k}))\>,$
where the approximation arises from the score function estimation and $\bm{x}_{k,T_{k}}$ is a sample at iteration $k$ diffused upto switching time $T_{k}$ via the forward process.
Compared to reverse SDE sampling, employing Tweedie’s formula will lead to a tradeoff between computation time and reconstruction quality. As evidenced by our experimental results in Section 5, using Tweedie’s formula for approximating the reverse SDE may result in blurry reconstructions with lower perceptual quality, a phenomenon partly attributed to the ”regression to the mean” effects [41].
Algorithm 2 LDM Version of Our Method

### v6

@@ 2.3 Diffusion Purification
Recently, diffusion purification has been shown to be an effective approach for removing adversarial perturbations [21, 29]. Specifically, for a given adversarially perturbed image $\bm{x}_{\textrm{adv}}$, diffusion purification first runs the forward process (2) starting from $\bm{x}_{\textrm{adv}}$ to some intermediate timestep $t^{\prime}$ such that $\bm{x}_{t^{\prime}}\sim p_{0t^{\prime}}(\bm{x}_{t^{\prime}}|\bm{x}_{\textrm{adv}})$:
$\bm{x}_{t^{\prime}}=\sqrt{\alpha_{t^{\prime}}}\bm{x}_{\textrm{adv}}+\sqrt{(1-\alpha_{t^{\prime}})}\bm{\epsilon}\>,$ (12)
@@
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Starting from $\bm{x}_{t^{\prime}}$, diffusion purification performs the reverse diffusion process (3) or (4) to produce a “purified” image $\bm{x}_{\textrm{purified}}$, effectively eliminating the adversarial perturbations. The rationale behind this approach is that by carefully selecting $t^{\prime}$, the adversarial noise within $\bm{x}_{\textrm{adv}}$ becomes overwhelmed by the noise introduced during the forward process, yet the core semantic and structural information is preserved. Consequently, executing the reverse diffusion process leads to a “clean” image reconstruction.
Distinct from existing methods that use diffusion purification to improve the robustness of DNN-based classifiers [21], in our work, we use diffusion purification as an operator that enforces the image prior embedded in the diffusion models.
3 Proposed Method
This section first motivates DCDP from the perspective of classical variable splitting techniques (section 3.1). Then we present our overall algorithm (section 3.2) and its extension to latent diffusion models and consistency models (sections 3.3 and 3.4).
3.1 Motivation: Methods with Variable Splitting
@@
Our method is inspired by the classical variable splitting techniques [30, 31, 32, 33] in optimization. Consider the following regularized optimization formulation for solving (1),
$\min_{\bm{x}}f(\bm{x})+\lambda\mathcal{R}(\bm{x}),$
where we typically adopt the least squares formulation, $f(\bm{x}):=\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$, as the data fidelity term to enforce consistency with the measurements, $\mathcal{R}(\cdot)$ is the regularizer that enforces the data prior, and $\lambda>0$ is a hyperparameter that balances the fidelity and the regularization strength. The idea behind variable splitting is to introduce an auxiliary variable $\bm{v}=\bm{x}$, resulting in an equivalent problem
$\min_{\bm{x},\bm{v}}f(\bm{x})+\lambda\mathcal{R}(\bm{v}),\quad\mathrm{s.t.}~{}~{}\;\bm{x}=\bm{v},$ (13)
@@
where the constrained optimization problem above can be further relaxed into an unconstrained problem by the half quadratic splitting technique (HQS) [34, 33]:
$\min_{\bm{x},\bm{v}}\mathcal{L}(\bm{x},\bm{v}):=f(\bm{x})+\lambda\,\mathcal{R}(\bm{v})+\mu\|\bm{x}-\bm{v}\|_{2}^{2},$ (14)
@@
for some properly chosen $\mu>0$. To solve (14), a common strategy is to adopt an iterative method that, at each iteration $k$, uses the auxiliary variable $\bm{v}$ to decompose the optimization into two subproblems of the form:
$\bm{x}_{k}=\arg\min_{\bm{x}}\;\mathcal{L}(\bm{x},\bm{v}_{k-1}),$ (15)
$\bm{v}_{k}=\arg\min_{\bm{v}}\;\mathcal{L}(\bm{x}_{k},\bm{v}).$ (16)
In this way, the data-consistency term $f(\bm{x})$ is decoupled from the regularization term $\mathcal{R}(\bm{v})$. Then, instead of having an explicit regularizer $\mathcal{R(\cdot)}$ (e.g., $\ell_{1}$-norm to induce sparsity), the regularization-based step (16) can be replaced with operators that implicitly enforce the data prior. In this work, we leverage diffusion purification to enforce the image prior encoded within the diffusion models as we discuss in Section 3.2.
3.2 Decoupled Data Consistency via Diffusion Purification
Our DCDP method alternates between the following two phases: (i) a reconstruction phase that enforces data consistency and (ii) a refinement phase that enforces the prior via diffusion purification. Below, we discuss these steps in more detail, with an illustration in Figure 1.
Reconstruction Phase. The first step of our algorithm is to solve the sub-problem (15) to enforce data consistency by minimizing the data fidelity term $f(\bm{x})$. Given $\bm{v}_{k-1}$ from the previous iteration, optimizing the loss $\mathcal{L}(\bm{x},\bm{v}_{k-1})$ for $\bm{x}$ depends only on the data fidelity term $f(\bm{x})=\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$ and its associated proximal term $\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}$, leading to:
$\bm{x}_{k}=\arg\min_{\bm{x}}\;\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}+\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}\>.$ (17)
Although (17) can be solved via iterative methods (e.g., SGD), selecting an appropriate hyperparameter $\mu$ often requires manual tuning. To mitigate this issue, we compute an approximate solution $\bm{x}_{k}$ by minimizing only the data consistency term $f(\bm{x})$, initialized at $\bm{v}_{k-1}$ and run for a fixed number of gradient steps. Empirically, this approach (i) yields a solution $\bm{x}_{k}$ that stays close to $\bm{v}_{k-1}$ without relying on $\mu$, and (ii) maintains low computational cost while being less sensitive to the hyperparameter.
Note that enforcing data consistency with (17) is significantly more efficient than DPS-based methods, as it does not require backpropagation over the entire score function. In practice, we can use any gradient-based optimizer (e.g., SGD, ADAM) for minimizing (17). Throughout the experiments, we fix the number of gradient steps $\tau$ for each round of the data consistency optimization. When the measurements are corrupted by noise, one can reduce $\tau$ to prevent overfitting. We discuss our choice of optimizers in detail in Appendix B. For simplicity, we denote this reconstruction phase by:
Refinement Phase. Due to the ill-posed nature of the problem in (1), $\bm{x}_{k}$ obtained in the reconstruction phase could suffer from task-specific artifacts, as shown in Figure 2 (third column). To alleviate these artifacts from $\bm{x}_{k}$, we propose to use diffusion purification described in Section 2.3 as an implicit regularizer. Specifically, let $t_{k}$ denote a predefined diffusion purification timestep that controls the strength of purification at iteration $k$. Following (12), we first define $\bm{x}_{k,0}:=\bm{x}_{k}$ and submerge this sample with noise by running the forward process up to time $t_{k}$:
$\bm{x}_{k,t_{k}}=\sqrt{\alpha_{t_{k}}}\bm{x}_{k,0}+\sqrt{1-\alpha_{t_{k}}}\bm{\epsilon}.$ (19)
Then, we run unconditional reverse sampling via either the reverse SDE (3) or the reverse ODE in (4) starting from the initial point $\bm{x}_{k,t_{k}}$ at the time $t_{k}$ to obtain the purified output $\bm{v}_{k}$. As demonstrated in Figure 2, diffusion purification effectively removes the artifacts in the reconstructions from the data consistency (fidelity) optimization (17) while preserving the overall image structures at the same time. For simplicity, we denote this refinement step by:
$\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},t_{k}).$ (20)
For the choice of $t_{k}$, we propose to use a linear decaying schedule for robust performance, which is capable of addressing the “hallucination” issue brought by the stochasticity of the diffusion purification process. In the early iterations of our method, when $t_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though it may be inconsistent with the measurements, will be free from artifacts and can serve as a good starting point for the subsequent iteration of data consistency optimization. As $t_{k}$ keeps decreasing when $k$ increases, the algorithm will converge to an equilibrium point which is both consistent with the measurements and free from the artifacts. The overall algorithm is presented in Algorithm 1. More technical details are deferred to Appendix D, and we empirically study the convergence in Section 4.3.
Algorithm 1 DCDP using Pixel Space Diffuison Models
@@ 3.2 Decoupled Data Consistency via Diffusion Purification
measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained diffusion model $\bm{s}_{\bm{\theta}}$, total number of reverse sampling steps $K$, decaying time schedule $\{t_{k}\}_{k\in[K]}$ for diffusion purification, the number of data consistency iterations $\tau$, and $\bm{v}_{0}=\bm{0}\in\mathbb{R}^{n}$
@@
     $\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau)$triangleright solve via gradient based methods for $\tau$ steps
@@
3.3 Fast Diffusion Purification for Refinement
In this subsection, we demonstrate two approaches to achieve faster diffusion purification: (i) by using accelerated samplers and (ii) by using Tweedie’s formula.
@@ 3.3 Fast Diffusion Purification for Refinement
Accelerated Samplers. To improve sampling efficiency, we can adapt accelerated samplers to speed up the reverse process of diffusion purification. For example, we can use DDIMs [8], which constructs a non-Markovian chain of the reverse process. By doing so, each round of diffusion purification can be performed with as few as $20$ unconditional reverse sampling steps (number of function evaluations, NFEs), with minimal performance degradation. We remark that existing methods such as DPS do not offer this advantage, as they typically require over $1000$ reverse sampling steps to achieve sufficient data consistency.
Approximation via Tweedie’s Formula. By Tweedie’s formula [35, 36], the posterior mean given noisy sample $\bm{x}_{t}$ can be estimated with the score function:
$\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})),$ (21)
@@
where $\bm{x}_{0}$ denotes a sample from the “clean” data distribution. Equipped with this formula, one can use the posterior mean as a single-step approximation to the solution of the reverse SDE (or ODE) in diffusion purification:
$\bm{v}_{k}$ $\approx\mathbb{E}\left[\bm{x}_{k,0}\mid\bm{x}_{k,t_{k}}\right]$ (22)
$\approx\frac{1}{\sqrt{\alpha_{t_{k}}}}(\bm{x}_{k,t_{k}}+(1-\alpha_{t_{k}})\bm{s}_{\bm{\theta}}(\bm{x}_{k,t_{k}},t_{k}))\>.$ (23)
Note that applying Tweedie’s formula leads to a trade-off between computation time and reconstruction quality. The multi-step reverse SDE produces a sharp sample that approximately follows the ground truth data distribution, whereas Tweedie’s formula returns the mean of all possible clean versions corresponding to the noisy sample $\bm{x}_{t}$, which is perceptually blurrier especially when $t$ is large. As evidenced by our experimental results in Section 4, using Tweedie’s formula may result in blurry reconstructions with lower perceptual quality. Nevertheless, DCDP with Tweedie’s formula still achieves competitive performance on standard metrics and is extremely fast as it requires only one step.
3.4 DCDP with Latent Diffusion and Consistency Models
By decoupling the data consistency steps from the diffusion sampling steps, our method enables the flexibility of utilizing a broad range of diffusion models in a plug-and-play fashion. To this end, we illustrate how to adapt DCDP to latent diffusion models and consistency models.
DCDP with Latent Diffusion Models (LDMs). LDMs [28] operate in a lower-dimensional latent space by training a pair of encoder and decoder [37] to improve the efficiency of vanilla diffusion models. Specifically, given an encoder $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ and a decoder $\mathcal{D}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ where $r\ll d$, LDMs reduce the computational complexity by mapping each $\bm{x}\in\mathbb{R}^{d}$ to $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{r}$ and performing the entire diffusion process in the latent space. After performing the reverse steps in the latent space, the sampled latent code $\bm{z}_{0}$ is decoded to a synthesized image via $\bm{x}_{0}=\mathcal{D}(\bm{z}_{0})$.
Unlike previous methods [18, 19], the flexibility of our framework allows us to (almost) seamlessly incorporate LDMs. For LDMs, we alternate between enforcing data consistency and diffusion purification, but both in the latent space:
$\widehat{\bm{v}}_{k}$ $=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},t_{k}),$ (25)
where $\mathcal{L}_{\mathrm{LDM}}$ is a loss function tailored for the latent space (see (26) and (29)), $\bm{s}_{\bm{\phi}}$ is the score function in the latent space, and $\bm{z}=\mathcal{E}(\bm{x})$ is the latent code. Once our method converges based on specific stopping criteria, we map the converged latent code back to the pixel space through the decoder $\mathcal{D}$. The overall method is presented in Algorithm 2. In the following subsections, we discuss two approaches for enforcing data consistency.
Approach I. Data Consistency in Latent Space
One approach for enforcing data consistency is to operate directly in the latent space by minimizing the following objective:
$\bm{z}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{z}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}+\mu\|\bm{z}-\widehat{\bm{v}}_{k-1}\|_{2}^{2}$ (26)
where $\widehat{\bm{v}}_{k-1}$ denotes the output from the previous diffusion purification step. We note that this optimization objective closely resembles the method in [4]. As discussed in Section 3.2, we approximately solve (26) by performing a limited number of gradient-based updates on the objective $\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$, initialized from $\widehat{\bm{v}}_{k-1}$ as:
Before diffusion purification, we run an extra step:
$\bm{z}_{k}$ $\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k})),$ (28)
which we denote by “Re-Encode”. The rationale behind Re-Encode is that due to the nonlinearity of the encoder $\mathcal{E}$, there could exist infinitely many solutions $\bm{z}$ for the optimization problem (27). However, since the LDMs are trained only on $\bm{z}=\mathcal{E}(\bm{x})$, where $\bm{x}$ follows the image distribution, we need to first map $\bm{z}_{k}$ back to the image space with $\bm{x}_{k}=\mathcal{D}(\bm{z}_{k})$ and then re-encode $\bm{x}_{k}$ as $\bm{z}_{k}=\mathcal{E}(\bm{x}_{k})$.
Algorithm 2 LDM Version of DCDP